(a) Define a law of electrostatics in integral form that is used to compute the electrostatic field E due to a symmetric distribution of charge within a given volume. State the meaning of the terms in the defining equation. (b) A uniformly charged long cylinder of radius a and length L has total charge q inside its volume. What is the direction of the electric field at points outside the cylinder? Find the electric field inside and outside the cylinder. (c) Total charge q is distributed along a line of length L. Determine the electric field produced by the wire using Gauss’s law. Compare this result with the result in part (b).


1
Expert's answer
2020-07-14T08:54:19-0400

A)

"\oint_S \vec E\vec{dS} = \iiint_V div(\vec E) dV" - Gauss theorem

"\iiint_V div(\vec E) dV = -4\pi\iiint_V \rho dV" - one of Maxwell equations

"-\frac{1}{4\pi}\oint_S \vec E\vec{dS} = \iiint_V \rho dV"

This equation says

The flow of the E-field through a surface S increase if and only if increase the total charge of volume V bonded by a surface S.

B) Cylinder (R - a radius, L - a length) case:

"\oint_S \vec E\vec{dS} = 2\pi LrE_r = -\frac{4\pi q}{\pi R^2 L}V_{cyl} = -4\pi q"

There is "E_r" - a radial component of the E - field

"E_r = -\frac{2q}{rL}", "r\leq R"

"E_r = -\frac{2q}{RL}, r> R"

C) Line cases(L - a length) case:

"\oint_S \vec E\vec{dS} = 2\pi LrE_r = -4\pi q"

"E_r = -\frac{2q}{rL}"

We assume q<0 in both cases